OpenStudy (anonymous):

find the maclaurin series in closed form of f(x)=1/sqrt(1-x)

OpenStudy (anonymous):

\[f(x)=(1-x)^{-1/2}\] expand this f(x)=1 + x/2 + 3/8(x^2) + ....